KCC Company has incurred the following costs on Job KC123, an order for 40 gearing wheels to be delivered at the end of next month. Direct materials: On March 5, requisition number 501 was issued for 40 titanium blanks to be used in the special order. The blanks cost P80.00 each. On March 8, requisition number 503 was issued for 960 hardened nibs also to be used in the special order. The nibs cost P6.00 each. Direct labor On March 9, Kristine worked from 9:00AM until 12:15PM on Job KC123. She is paid P120.00 per hour. On March 21, Clarisse worked from 2:15PM until 4:30PM on Job KC123. She is paid P140.00 per hour. Required: 1. On what documents would these costs be recorded? 2. How much cost should have been recorded on each of the documents for Job KC123?
Answer #1

1)Direct materials costs are recorded in the material requisition form and direct labor costs are recorded in the Employee time ticket
2)In material requisition form the direct material costs is recorded and total cost:
=(40*80)+(960*6)
=8960

In employee time ticket the direct labour costs are recorded and total costs is beloe
Kristine total hours worked=9am to 12.15pm= 3hr 15min and cost is 120 per hour
=(120*3)+(0.25*120)=390
Clarrise total hrs worked=2.15pm to 4.30pm=2hr 15min
=(120*2)+(0.25*120)=270
=390+270=660